### v1

@@ 1 Introduction
Two technical challenges arise in conducting uniform inference. First, the (asymptotic) joint distribution of $M_{i,T}$ for $i=1,\ldots,d$ depends on unknown parameters, making it non-pivotal. To address this, we propose a resampling method based on the moving blocks bootstrap (MBB) (Kunsch, 1989; Liu and Singh, 1992; Fitzenberger, 1998) to approximate the finite-sample “null” joint distribution of $M_{i,T}$ for $i=1,\ldots,d$. The MBB also accounts for potential serial correlation in financial returns. Second, $M_{i,T}$ diverges to infinity with $T$ if and only if $i\in\mathrm{MSS}$. To ensure the inference procedure is not conservative and maintains exact control of size (Type-I error), the desired “null” joint distribution of $M_{i,T}$ should be concentrated on $i\notin\mathrm{MSS}$. However, since the $\mathrm{MSS}$ is unknown, this desired “null” joint distribution remains infeasible even with the MBB. We address this issue through a two-step approach. In the first-step, we compute a critical value based on a known upper bound of the $\mathrm{MSS}$. While conservative, this critical value ensures consistent estimation of the $\mathrm{MSS}$. In the second-step, the consistent $\mathrm{MSS}$ estimator obtained in the first-step is used to refine the critical value, enabling non-conservative and more powerful inference on the $\mathrm{MSS}$.
@@
Our estimator of the $\mathrm{MSS}$ is formally defined as the subset of assets whose $M_{i,T}$ exceeds the refined critical value obtained through the two-step approach. Additionally, the magnitude of $M_{i,T}$ serves as a metric for evaluating the relative importance of the assets and ranking them in $R$. We theoretically demonstrate that this $\mathrm{MSS}$ estimator covers the true $\mathrm{MSS}$ with probability approaching 1 (wpa1), and converges to the exact $\mathrm{MSS}$ with probability reaching any pre-specified level, such as 0.95 or 0.99. This estimator can be made consistent by lettng the pre-specified level approach 1 with increasing sample size.
@@
As a by-product, our $\mathrm{MSS}$ estimation procedure can also be applied to the conventional spanning test problem, offering more insights than traditional spanning tests. Given a pre-specified benchmark asset set and an additional asset set, our method can identify and estimate the $\mathrm{MSS}$ of all assets under consideration. To determine whether the benchmark set spans the additional set, we simply check whether the estimated $\mathrm{MSS}$ is a subset of the benchmark set. More importantly, by analyzing the intersections of the estimated $\mathrm{MSS}$ with the benchmark set and the additional set, we can identify which assets in the benchmark set become redundant upon including the additional set, and which assets in the additional set are truly valuable. This approach offers a more nuanced and refined assessment of asset relevance, surpassing the binary conclusions of conventional spanning tests.
@@
The finite-sample performance of our proposed $\mathrm{MSS}$ estimation procedure is assessed through extensive Monte Carlo simulations. We simulate the data using a model with an autoregressive (AR) conditional mean and a generalized autoregressive conditional heteroskedasticity (GARCH) conditional variance, which effectively captures key stylized features of financial returns, including serial and cross-sectional correlation as well as volatility clustering. The simulation results demonstrate that the empirical probability of the estimated $\mathrm{MSS}$ containing the true $\mathrm{MSS}$ approaches one as the sample size increases. Furthermore, the empirical probability of the estimated $\mathrm{MSS}$ being identical to the true $\mathrm{MSS}$ aligns closely with the nominal significance level for sufficiently large sample sizes. These findings are consistent with the asymptotic theory estalished for our method, demonstrating its robust performance in finite samples.
@@
Our study makes a direct contribution to the growing literature on conventional spanning tests. For example, Huberman and Kandel (1987) derives the key conditions under which a given set of assets spans the mean-variance frontier of a larger set when additional assets are included, and introduces a likelihood ratio test to assess the redundancy of the additional set of assets. Subsequent advancements in this field have been made by Hansen and Jagannathan (1991), Ferson, Foerster, and Keim (1993), De Santis (1993), Bekaert and Urias (1996), De Roon, Nijman, and Werker (2001), Amengual and Sentana (2010), Kan and Zhou (2012), Peñaranda and Sentana (2012), among others. However, as emphasized earlier, our study is the first to focus on the identification and estimation of the $\mathrm{MSS}$, marking a significant departure from the existing literature on spanning tests. Empirically, our work adds to the ongoing discussions on the interplay between factor momentum and momentum factors, as detailed by Gupta and Kelly (2018), Ehsani and Linnainmaa (2022), Yan and Yu (2023), and Arnott, Kalesnik, and Linnainmaa (2023). By characterizing the $\mathrm{MSS}$ for various momentum strategies, evaluating their relative importance, and ranking them within a large set of assets, our approach provides novel insights into the interactions between these factors and their role in mean-variance analysis.
@@
The remainder of this paper is organized as follows. Section 2 introduces the identification conditions for the MSS and details the implementation of the proposed estimation and inference method. This section also establishes the asymptotic properties of the method. Section 3 presents simulation studies to assess the finite-sample performance of the method, while Section 4 offers an empirical application. Finally, Section 5 concludes the paper. The Appendix includes proofs of the main theoretical results, auxiliary lemmas used in these proofs, and additional simulation results. The Supplemental Appendix contains detailed proofs of the auxiliary lemmas.
@@
Notation. We use $a\equiv b$ to indicate that $a$ is defined as $b$. For any positive integer $m$, let $I_{m}$ denote the $m\times m$ identity matrix. For any positive integers $m_{1}$ and $m_{2}$, $\mathbf{1}_{m_{1}\times m_{2}}$ and $\mathbf{0}_{m_{1}\times m_{2}}$ denote the $m_{1}\times m_{2}$ matrices of ones and zeros, respectively. For real numbers $a_{1},\ldots,a_{m}$, let $(a_{i})_{i\leq m}\equiv(a_{1},\ldots,a_{m})^{\top}$, and let $a_{-i}$ denote the subvector of $(a_{i})_{i\leq m}$ with $a_{i}$ excluded. Define the support of $(a_{i})_{i\leq m}$ as $\mathrm{Supp}_{(a_{i})_{i\leq m}}\equiv\{i=1,\ldots,m:a_{i}\neq 0\}$. For any matrices $A$ and $B$, $\mathrm{diag}(A,B)$ represents a block diagonal matrix with $A$ and $B$ as its diagonal blocks, and $A\otimes B$ denotes the Kronecker product of $A$ and $B$. Additionally, $A_{j,.}$ represents the $j$th row of the matrix $A$. For any positive integer $d$, let $\mathcal{M}\equiv\{1,\ldots,d\}$, and for any positive integer $i\leq d$, let $\ell_{d,i}$ denote the $d\times 1$ vector whose $i$-th entry is $1$, with all other entries equal to 00. For two sequences of positive numbers $a_{n}$ and $b_{n}$, we write $a_{n}\succ b_{n}$ if $a_{n}\geq c_{n}b_{n}$ for some strictly positive sequence $c_{n}\rightarrow\infty$.
@@ 2 Main Theory
This section presents the main theoretical results of the paper. In subsection 2.1, we examine the existence and uniqueness of the $\mathrm{MSS}$ and establish its identification condition. The identification condition is constructive, as it is employed in subsection 2.2 to develop valid estimation and inference methods for the $\mathrm{MSS}$.
@@ 2.1 Identification Condition
Since the mean-variance frontier of $R$ depends only on its mean $\mu$ and variance $\Sigma$, which are consistently estimable, we assume in this subsection that both $\mu$ and $\Sigma$ are known for the purpose of investigating the identification condition of the $\mathrm{MSS}$. Given that the set of assets $R$ spans its own mean-variance frontier, the $\mathrm{MSS}$ is guaranteed to exist. Specifically, since there are $22^{d}-$ 2 nonempty and proper subvectors of $R$, we can examine each of these subvectors and identify those that span the mean-variance frontier of $R$. The $\mathrm{MSS}$ will be the subvector(s) in this collection with the smallest dimension. If no proper subvector spans the mean-variance frontier of $R$, then the $\mathrm{MSS}$ is $R$ itself.
@@
$\mathrm{MSS}=\left\{i=1,\ldots,d\text{: }\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}>0\right\}.$ (2.3)
Second, the identification condition presented in Lemma 1 is constructive and does not require any prior knowledge of the $\mathrm{MSS}$. In the population where $\mu$ and $\Sigma$ of $R$ are known, we only need to compute $\theta_{i}$ for $i=1,\ldots,d$, and use (2.3) to determine the $\mathrm{MSS}$. Third, Lemma 1 strengthens a key result from Huberman and Kandel (1987), which is widely used in practice to assess the redundancy of additional assets relative to a benchmark asset set. Specifically, Proposition 3 of Huberman and Kandel (1987) implies that $R_{-i}$ spans the mean-variance frontier of $R$ (and hence $R_{i}$ is redundant relative to $R_{-i}$) if and only if:
@@
In practice, the values of $\theta_{i}$ for each asset $R_{i}$ are unknown but can be estimated through LS regression of $R_{i}$ on $R_{-i}$. Combined with the identification conditions established in Lemma 1, this enables the estimation and inference of the $\mathrm{MSS}$ in finite samples. In the next subsection, we analyze the asymptotic properties of the LS estimators of $\theta_{i}$ uniformly over $i=1,\ldots,d$. These properties facilitate the construction of $\mathrm{MSS}$ estimators that cover the true $\mathrm{MSS}$ wpa1, and asymptotically identify it with any desired level of confidence.
@@ 2.2 Estimation and Inference of the Minimum Spanning Set
We first introduce some notations to simplify the definition of the estimator for $\theta_{i}$, as well as the estimation procedure for the $\mathrm{MSS}$. The key value $\theta_{i}$ can be expressed as a linear transformation of $\alpha_{i}$ and $\beta_{i}$: $\theta_{i}=A\cdot(\alpha_{i},\beta_{i}^{\top})^{\top}$ where $A\equiv\mathrm{diag}(1,1_{d-1}^{\top})$. For any observation $R_{t}\equiv(R_{i,t})_{i\leq d}$ where $t=1,\ldots T$, we let $\tilde{R}_{-i,t}\equiv[1,R_{-i,t}^{\top}]^{\top}$ denote the regressors in the linear regression specified in (2.1), $\hat{Q}_{-i}\equiv T^{-1}\sum_{t\leq T}\tilde{R}_{-i,t}\tilde{R}_{-i,t}^{\top}$ and $Q_{-i}\equiv\mathrm{E}[\tilde{R}_{-i,t}\tilde{R}_{-i,t}^{\top}]$ for any $i=1,\ldots,d$.
Next, we describe the estimation procedures for the $\mathrm{MSS}$ and their intuition. For each asset $R_{i}$, the LS estimators of $\alpha_{i}$ and $\beta_{i}$ are defined as:
$(\hat{\alpha}_{i},\hat{\beta}_{i}^{\top})^{\top}\equiv\hat{Q}_{-i}T^{-1}\sum_{t\leq T}\tilde{R}_{-i,t}R_{i,t}.$ (2.5)
@@
This pointwise result can be used to test whether a given asset $R_{i}$ belongs to the $\mathrm{MSS}$ or not. However, our goal is to estimate and conduct statistical inference on the $\mathrm{MSS}$, which may include multiple assets. Therefore, to ensure accurate estimation of the $\mathrm{MSS}$ and proper control of statistical inference errors, we need to conduct a joint statistical inference on condition (2.2) for $i=1,\ldots,d$, which requires approximating the finite-sample distribution of $\hat{\theta}_{i}$ uniformly over $i=1,\ldots,d$.
@@
$\tilde{M}_{i,T}\equiv\left\|(\ell_{d,i}^{\top}\otimes\mathrm{diag}(s_{i,1}^{-1},s_{i,2}^{-1}))\Omega_{d,T}^{1/2}\mathcal{N}_{d}\right\|_{\infty},$ (2.10)
where $s_{i,j}^{2}\equiv\sigma_{\varepsilon_{i}}^{2}A_{j,\cdot}Q_{-i}^{-1}A_{j,\cdot}^{\top}$ and $\sigma_{\varepsilon_{i}}^{2}\equiv\mathrm{E}[\varepsilon_{i,t}^{2}]$. This implies that the assets in the $\mathrm{MSS}$ tend to have larger $M_{i,T}$ values than those that are not in the $\mathrm{MSS}$. A formal statistical inference procedure should provide a critical value under a pre-specified significance level $p\in(0,1)$ to determine when an asset with a large $M_{i,T}$ value can be included in the $\mathrm{MSS}$.
Ideally, for any small $p\in(0,1)$, the critical value should depend only on the assets that are not in the $\mathrm{MSS}$. This is because, in view of Lemma 1 the “null hypothesis”:
@@
holds only for $i\notin\mathrm{MSS}$. Therefore, if the $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$, we would use the $(1-p)$-quantile of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, denoted as $cv_{1-p}^{u}$, as the critical value. On the other hand, if $\mathrm{MSS}=\mathcal{M}$, we shall set $cv_{1-p}^{u}$ to any fixed constant. From the discussion in the previous paragraph, as $T$approaches infinity, we have
$\mathrm{P}\left(\min_{i\in\mathrm{MSS}}M_{i,T}>cv_{1-p}^{u}\right)\rightarrow 1$ (2.11)
regardless of whether the $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$, and
$\mathrm{P}\left(\max_{i\notin\mathrm{MSS}}M_{i,T}>cv_{1-p}^{u}\right)\rightarrow p,$ (2.12)
if $\mathrm{MSS}$ is a proper subset of $\mathcal{M}$. This implies that the estimator of $\mathrm{MSS}$ defined as:
$\widehat{\mathrm{MSS}}_{p}^{u}\equiv\left\{i=1,\ldots,d\text{: }M_{i,T}>cv_{1-p}^{u}\right\}$ (2.13)
has the desirable properties that it covers the true $\mathrm{MSS}$ wpa1, and overestimates it with a small probability $p$.
However, the critical value $cv_{1-p}^{u}$ is not practical to use for two main reasons: (i) the distribution of $(\tilde{M}_{i,T})_{i\leq d}$ is unknown due to the presence of nuisance parameters $(s_{i,j})_{j\leq 2}$ ($i\leq d$) and $\Omega_{d,T}$; and (ii) there is no prior knowledge about the $\mathrm{MSS}$. The first challenge can be addressed by estimating the distribution of $(\tilde{M}_{i,T})_{i\leq d}$ through either plugging in consistent estimators of the nuisance parameters $(s_{i,j})_{j\leq 2}$ ($i\leq d$) and $\Omega_{d,T}$ directly, or by using resampling methods such as the bootstrap. In this paper, we employ the moving blocks (Kunsch, 1989; Liu and Singh, 1992; Fitzenberger, 1998) bootstrap (MBB) to approximate the distribution of $(\tilde{M}_{i,T})_{i\leq d}$.
The second challenge is more delicate. One solution is to use a known upper bound of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, e.g., $\max_{i\leq d}\tilde{M}_{i,T}$, to obtain a critical value which is larger than $cv_{1-p}^{u}$. While straightforward to implement, this method is conservative, potentially leading to an underestimation of the $\mathrm{MSS}$ in finite samples. An alternative approach involves replacing the $\mathrm{MSS}$ in $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$ with a preliminary consistent estimator, which can be obtained using the first method by letting $p$ approach zero. In this paper, we study both methods and refer to their corresponding $\mathrm{MSS}$ estimators as the one-step estimator and the two-step estimator, respectively.
Let $\hat{\alpha}_{i}^{b}$, $\hat{\beta}_{i}^{b}$ and $\hat{s}_{i,j}^{b}$ denote the bootstrap counterparts of $\hat{\alpha}_{i}$, $\hat{\beta}_{i}$ and $\hat{s}_{i,j}$, respectively. [^2] The details of the MBB bootstrap are provided in the algorithm below. For $i\leq d$, define:

### v3

@@ 1 Introduction
Two technical challenges arise in conducting uniform inference. First, the (asymptotic) joint distribution of $M_{i,T}$ for $i=1,\ldots,d$ depends on unknown parameters, making it non-pivotal. To address this, we propose a resampling method based on the moving blocks bootstrap (MBB) (Kunsch, 1989; Liu and Singh, 1992; Fitzenberger, 1998) to approximate the finite-sample “null” joint distribution of $M_{i,T}$ for $i=1,\ldots,d$. The MBB also accounts for potential serial correlation in financial returns. Second, $M_{i,T}$ diverges to infinity with $T$ if and only if $i\in\mathrm{MSS}$. To ensure the inference procedure is not conservative and maintains exact control of size (Type-I error), the desired “null” joint distribution of $M_{i,T}$ should be concentrated on $i\notin\mathrm{MSS}$. However, since the $\mathrm{MSS}$ is unknown, this desired “null” joint distribution remains infeasible even with the MBB. To address this issue, we adopt a step-down approach from the multiple hypothesis testing literature (see, e.g., Romano and Wolf (2005)). This method iteratively refines the bootstrap critical value used for $\mathrm{MSS}$ estimation, improving the power of our procedure in identifying important assets in $R$.
@@
Our estimator of the $\mathrm{MSS}$ is formally defined as the subset of assets whose $M_{i,T}$ exceeds the refined bootstrap critical value obtained through the step-down approach. Additionally, the magnitude of $M_{i,T}$ serves as a metric for evaluating the relative importance of the assets and ranking them in $R$. We theoretically demonstrate that this $\mathrm{MSS}$ estimator covers the true $\mathrm{MSS}$ with probability approaching 1 (wpa1), and converges to the exact $\mathrm{MSS}$ with probability reaching any pre-specified level, such as 0.95 or 0.99. Additionally, this estimator can be made consistent by letting the pre-specified level approach 1 with increasing sample size.
@@
As a by-product, our $\mathrm{MSS}$ estimation procedure can also be applied to the conventional spanning test problem, offering more insights than traditional spanning tests. Given a pre-specified benchmark asset set and an additional asset set, our method can identify and estimate the $\mathrm{MSS}$ of all assets under consideration. To determine whether the benchmark set spans the additional set, we can simply check whether the estimated $\mathrm{MSS}$ is a subset of the benchmark set. More importantly, by analyzing the intersections of the estimated $\mathrm{MSS}$ with the benchmark set and the additional set, we can identify which assets in the benchmark set become redundant upon including the additional set, and which assets in the additional set are truly valuable. This approach provides a more nuanced and refined assessment of asset relevance, surpassing the binary conclusions of conventional spanning tests.
@@
The finite-sample performance of our proposed $\mathrm{MSS}$ estimation procedure is assessed through extensive Monte Carlo simulations. We simulate the data using a model with an autoregressive (AR) conditional mean and a generalized autoregressive conditional heteroskedasticity (GARCH) conditional variance, which effectively captures key stylized features of financial returns, including serial and cross-sectional correlation as well as volatility clustering. The simulation results demonstrate that the empirical probability of the estimated $\mathrm{MSS}$ containing the true $\mathrm{MSS}$ approaches one as the sample size increases. Furthermore, the empirical probability of the estimated $\mathrm{MSS}$ being identical to the true $\mathrm{MSS}$ aligns closely with the nominal significance level for sufficiently large sample sizes. These findings are consistent with the asymptotic theory established for our method, demonstrating its robust performance in finite samples.
@@
Our study makes a direct contribution to the growing literature on conventional spanning tests. Huberman and Kandel (1987) derives the key conditions under which a given set of assets spans the mean-variance frontier of a larger set when additional assets are included, and introduces a likelihood ratio test to assess the redundancy of the additional set of assets. Subsequent advancements in this field have been made by Hansen and Jagannathan (1991), Ferson, Foerster, and Keim (1993), De Santis (1993), Bekaert and Urias (1996), De Roon, Nijman, and Werker (2001), Amengual and Sentana (2010), Kan and Zhou (2012), Peñaranda and Sentana (2012), among others. However, as emphasized earlier, our study is the first to focus on the identification and estimation of the $\mathrm{MSS}$, marking a significant departure from the existing literature on spanning tests. Empirically, our work adds to the ongoing discussions on the interplay between factor momentum and momentum factors, as detailed by Gupta and Kelly (2018), Ehsani and Linnainmaa (2022), Yan and Yu (2023), and Arnott, Kalesnik, and Linnainmaa (2023). By characterizing the $\mathrm{MSS}$ for various momentum strategies, evaluating their relative importance, and ranking them within a large set of assets, our approach provides novel insights into the interactions between these factors and their role in mean-variance analysis.
@@
The remainder of this paper is organized as follows. Section 2 introduces the identification conditions for the MSS and details the implementation of the proposed estimation and inference method. It also establishes the method’s asymptotic properties and demonstrates how slight modifications can extend its applicability to other important problems in empirical finance. Section 3 presents simulation studies to assess the finite-sample performance of the method, while Section 4 offers an empirical application. Finally, Section 5 concludes the paper. The Appendix includes proofs of the main theoretical results, auxiliary lemmas used in these proofs, and additional simulation results. The Supplemental Appendix contains detailed proofs of the auxiliary lemmas.
@@
Notation. We use $a\equiv b$ to indicate that $a$ is defined as $b$. For any positive integer $m$, let $I_{m}$ denote the $m\times m$ identity matrix. For any positive integers $m_{1}$ and $m_{2}$, $\mathbf{1}_{m_{1}\times m_{2}}$ and $\mathbf{0}_{m_{1}\times m_{2}}$ denote the $m_{1}\times m_{2}$ matrices of ones and zeros, respectively. For real numbers $a_{1},\ldots,a_{m}$, let $(a_{i})_{i\leq m}\equiv(a_{1},\ldots,a_{m})^{\top}$, and let $a_{-i}$ denote the subvector of $(a_{i})_{i\leq m}$ with $a_{i}$ excluded. The $\ell_{\infty}$ norm of $(a_{i})_{i\leq m}$ is given by $|(a_{i})_{i\leq m}|_{\infty}=\max_{i\leq m}\left|a_{i}\right|$. Define the support of $(a_{i})_{i\leq m}$ as $\mathrm{Supp}_{(a_{i})_{i\leq m}}\equiv\{i=1,\ldots,m:a_{i}\neq 0\}$. For any matrices $A$ and $B$, $\mathrm{diag}(A,B)$ represents a block diagonal matrix with $A$ and $B$ as its diagonal blocks, and $A\otimes B$ denotes the Kronecker product of $A$ and $B$. Additionally, $A_{j,.}$ represents the $j$th row of the matrix $A$. For any positive integer $d$, let $\mathcal{M}_{d}\equiv\{1,\ldots,d\}$, and for any positive integer $i\leq d$, let $\ell_{d,i}$ denote the $d\times 1$ vector whose $i$th entry is $1$, with all other entries equal to 00. For two sequences of positive numbers $a_{n}$ and $b_{n}$, we write $a_{n}\succ b_{n}$ if $a_{n}\geq c_{n}b_{n}$ for some strictly positive sequence $c_{n}\rightarrow\infty$.
@@ 2 Main Theory
This section presents the main theoretical results of the paper. In subsection 2.1, we examine the existence and uniqueness of the $\mathrm{MSS}$ and establish its identification condition. The identification condition is constructive, as it is employed in subsection 2.2 to develop valid estimation and inference method for the $\mathrm{MSS}$. Subsection  2.3 demonstrates how the $\mathrm{MSS}$ estimation and inference procedure can be adapted for other important applications in empirical finance.
@@ 2.1 Identification Condition
Since the mean-variance frontier of $R$ depends only on its mean $\mu$ and variance $\Sigma$, which are consistently estimable, we assume in this subsection that both $\mu$ and $\Sigma$ are known for the purpose of investigating the identification condition of the $\mathrm{MSS}$. Given that the set of assets $R$ spans its own mean-variance frontier, the $\mathrm{MSS}$ is guaranteed to exist. Specifically, since there are $22^{d}-$ 2 nonempty and proper subvectors of $R$, we can examine each of these subvectors and identify those that span the mean-variance frontier of $R$. The $\mathrm{MSS}$ will be the subvector(s) in this collection with the smallest dimension. If no proper subvector of $R$ spans the mean-variance frontier of $R$, then the $\mathrm{MSS}$ is $R$ itself.
@@
$\mathrm{MSS}=\left\{i\in\mathcal{M}_{d}\text{: }\alpha_{i}^{2}+(1_{d-1}^{\top}\beta_{i}-1)^{2}>0\right\}.$ (2.3)
Second, the identification condition presented in Lemma 1 is constructive and does not require any prior knowledge of the $\mathrm{MSS}$. Given the population values $\mu$ and $\Sigma$, we only need to compute $\theta_{i}$ for $i\in\mathcal{M}_{d}$, and use (2.3) to determine the $\mathrm{MSS}$. Third, Lemma 1 strengthens a key result from Huberman and Kandel (1987), which is widely used in the literature to assess the redundancy of additional assets relative to a benchmark asset set. Specifically, Proposition 3 of Huberman and Kandel (1987) implies that $R_{-i}$ spans the mean-variance frontier of $R$ (and hence $R_{i}$ is redundant relative to $R_{-i}$) if and only if:
@@
In practice, the value of $\theta_{i}$ for each asset $R_{i}$ is unknown but can be estimated through LS regression of $R_{i}$ on $R_{-i}$. Combined with the identification conditions established in Lemma 1, this enables the estimation and inference of the $\mathrm{MSS}$ in finite samples. In the next subsection, we analyze the asymptotic properties of the LS estimators of $\theta_{i}$ uniformly over $i\in\mathcal{M}_{d}$. These properties facilitate the construction of $\mathrm{MSS}$ estimator that cover the true $\mathrm{MSS}$ with probability approaching 1 (wpa1), and asymptotically identify it with any desired level of confidence.
@@ 2.2 Estimation and Inference of the Minimum Spanning Set
We first introduce some notations to simplify the definition of the estimator for $\theta_{i}$, as well as the estimation procedure for the $\mathrm{MSS}$. The key value $\theta_{i}$ can be expressed as a linear transformation of $\alpha_{i}$ and $\beta_{i}$: $\theta_{i}\equiv A\cdot(\alpha_{i},\beta_{i}^{\top})^{\top}$ where $A\equiv\mathrm{diag}(1,1_{d-1}^{\top})$. For any observation $R_{t}\equiv(R_{i,t})_{i\leq d}$ where $t=1,\ldots T$, we let $\tilde{R}_{-i,t}\equiv[1,R_{-i,t}^{\top}]^{\top}$ denote the regressors in the linear regression specified in (2.1), $\hat{Q}_{-i}\equiv T^{-1}\sum_{t\leq T}\tilde{R}_{-i,t}\tilde{R}_{-i,t}^{\top}$ and $Q_{-i}\equiv\mathrm{E}[\tilde{R}_{-i,t}\tilde{R}_{-i,t}^{\top}]$ for any $i\in\mathcal{M}_{d}$.
Next, we describe the estimation procedures for the $\mathrm{MSS}$ and their intuition. For each asset $R_{i}$, the LS estimator of $(\alpha_{i},\beta_{i}^{\top})^{\top}$ is defined as:
$(\hat{\alpha}_{i},\hat{\beta}_{i}^{\top})^{\top}\equiv\hat{Q}_{-i}^{-1}T^{-1}\sum_{t\leq T}\tilde{R}_{-i,t}R_{i,t}.$ (2.5)
@@
This pointwise result can be used to test whether a given asset $R_{i}$ belongs to the $\mathrm{MSS}$ or not. However, our goal is to estimate and conduct statistical inference on the $\mathrm{MSS}$, which may include multiple assets. Therefore, to ensure accurate estimation of the $\mathrm{MSS}$ and proper control of statistical inference errors, we need to conduct a joint statistical inference on condition (2.2) for $i\in\mathcal{M}_{d}$, which requires approximating the finite-sample distribution of $\hat{\theta}_{i}$ uniformly over $i\in\mathcal{M}_{d}$.
@@
$\tilde{M}_{i,T}\equiv\left|(\ell_{d,i}^{\top}\otimes\mathrm{diag}(s_{i,1}^{-1},s_{i,2}^{-1}))\Omega_{d,T}^{1/2}\mathcal{N}_{d}\right|_{\infty},$ (2.10)
where $s_{i,j}^{2}\equiv\sigma_{\varepsilon_{i}}^{2}A_{j,\cdot}Q_{-i}^{-1}A_{j,\cdot}^{\top}$ and $\sigma_{\varepsilon_{i}}^{2}\equiv\mathrm{E}[\varepsilon_{i,t}^{2}]$. Since the assets in the $\mathrm{MSS}$ tend to have larger $M_{i,T}$ values than those that are not in the $\mathrm{MSS}$, a formal statistical inference procedure should provide a critical value under a pre-specified significance level $p\in(0,1)$ to determine when an asset with a large $M_{i,T}$ value can be included in the $\mathrm{MSS}$.
Ideally, for any small $p\in(0,1)$, the critical value should depend only on the assets that are not in the $\mathrm{MSS}$. This is because, in view of Lemma 1, the  “null hypothesis”:
@@
holds only for $i\notin\mathrm{MSS}$. Therefore, if the $\mathrm{MSS}$ is a proper subset of $\mathcal{M}_{d}$, we would use the $(1-p)$-quantile of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, denoted as $cv_{1-p}^{u}$, as the critical value. This leads to an infeasible estimator of $\mathrm{MSS}$:
$\widehat{\mathrm{MSS}}_{p}^{u}\equiv\left\{i\in\mathcal{M}_{d}\text{: }M_{i,T}>cv_{1-p}^{u}\right\}.$ (2.11)
Under certain mild conditions, such as the sufficient conditions stated in Theorem 1 below, $\widehat{\mathrm{MSS}}_{p}^{u}$ possesses the desirable properties of covering the true $\mathrm{MSS}$ wpa1, and overestimates it with a small probability $p$.
However, the critical value $cv_{1-p}^{u}$ is not practical to use for two main reasons: (i) the distribution of $(\tilde{M}_{i,T})_{i\leq d}$ is unknown due to the presence of nuisance parameters $(s_{i,j})_{j\leq 2}$ ($i\leq d$) and $\Omega_{d,T}$; and (ii) there is no prior knowledge about the $\mathrm{MSS}$. The first challenge can be addressed by estimating the distribution of $(\tilde{M}_{i,T})_{i\leq d}$ through either plugging in consistent estimators of the nuisance parameters $(s_{i,j})_{j\leq 2}$ ($i\leq d$) and $\Omega_{d,T}$ directly, or by using resampling methods such as the bootstrap. In this paper, we employ the moving blocks (Kunsch, 1989; Liu and Singh, 1992; Fitzenberger, 1998) bootstrap (MBB) to approximate the distribution of $(\tilde{M}_{i,T})_{i\leq d}$. The second challenge is more delicate. One solution is to use a known upper bound of $\max_{i\notin\mathrm{MSS}}\tilde{M}_{i,T}$, e.g., $\max_{i\leq d}\tilde{M}_{i,T}$, to obtain a critical value which is larger than $cv_{1-p}^{u}$. While straightforward to implement, this method is conservative, potentially leading to an underestimation of the $\mathrm{MSS}$ in finite samples. Instead, we adopt the step-down procedure (see, e.g., Romano and Wolf (2005)), which iteratively refines the bootstrap critical value, thereby enhancing the power of our procedure in estimating the $\mathrm{MSS}$.
Details of our $\mathrm{MSS}$ estimation procedure are provided in the algorithm below.
Algorithm: A Bootstrap $\mathbf{MSS}$ Estimation Procedure
Step 3. Let $m\equiv[T/\ell]$ and $T_{B}\equiv m\ell$. Resample $m$ blocks $Z_{j}^{b}\equiv\{R_{\ell(j-1)+1}^{b},\ldots,R_{\ell(j-1)+\ell}^{b}\}$ independently from $\{B_{j}\}_{j\leq q}$ to obtain a bootstrap sample $\{R_{t}^{b}\}_{t\leq T_{B}}\equiv\{Z_{j}^{b}\}_{j\leq m}$, and then calculate